Pong Incorporated's segmented income statement for the most recent month is given below.
Total Store A Store B
  Sales $159,100 $65,700 $93,400
  Variable expenses 59,279 32,193 27,086
  Contribution margin 99,821 33,507 66,314
  Traceable fixed expenses 69,100 20,800 48,300
  Segment margin 30,721 $12,707 $18,014
  Common fixed expenses 24,800
  Net operating income $ 5,921

If Store B sales increase by $47,400 with no change in fixed expenses, the overall company net operating income should:
Answer #1

Store B Contribution margin ratio = 66314/93400 = 71%

Net operating income increase by = 47400*71% = 33654

If Store B sales increase by $47,400 with no change in fixed expenses, the overall company net operating income should: $33654